# PIRB: A Comprehensive Benchmark of Polish Dense and Hybrid Text Retrieval Methods

### 20 Feb 2024 v1

@@ 4.2.   Web datasets
The PIRB benchmark includes nine datasets crawled from Polish websites. Our goal was to enrich the benchmark with real-world data written originally in Polish. To accomplish this, we identified websites with separate Q&A sections, some containing questions and answers written directly by the site’s editors, and others allowing content creation by registered users. We selected a diverse set of data sources covering various domains such as medicine, law, business, physics, and linguistics. For each data source, we implemented a separate parser to enable precise extraction of question and answer content. We also conducted data cleaning and anonymization, removing personal and contact information. A summary of statistics regarding the collected datasets is presented in Table 1.
@@ 4.3.   GPT-exams
    In the last step, we manually removed from the dataset the cases in which the model refused to answer the question. We searched for phrases such as "model języka" (language model), "nie jestem" (I’m not), or "nie mogę" (I can’t).
@@ 5.1.   Baseline methods
  • BM25 (Robertson et al., 2009) is a popular and effective term-based ranking function used in information retrieval since the 1980s. It is an extension of TF-IDF weighting scheme, providing a balanced approach to text ranking, considering both term frequency and document specificity. In our experiments, we used the implementation of BM25 available in anserini (Yang et al., 2018).
  • Sparse Lexical and Expansion (SPLADE) model (Formal et al., 2021b, a) is a family of modern term-based methods employing Transformer language models. In this approach, the masked language modeling (MLM) head is optimized to generate a vocabulary-sized weight vector adapted for text retrieval. SPLADE is a highly effective sparse retrieval ranking algorithm, achieving better performance than classic methods such as BM25. We trained the Polish version following the same procedure as the SPLADE++ variant (Formal et al., 2022), utilizing hard negatives mined with an ensemble of cross-encoders (EnsembleDistil). We used Polish DistilRoBERTa (Dadas et al., 2020) as our base language model.
@@ 5.2.   Other methods
As part of our experiments, we evaluated over 20 publicly available Polish and multilingual dense text encoders, but only a few of them proved to be effective enough for practical use in retrieval tasks. In this publication, we present results only for the models that achieved an NDCG@10 score higher than the BM25 baseline. In the full ranking, which we have made available online555huggingface.co/spaces/sdadas/pirb, we include all of the evaluated retrievers, along with more detailed results covering individual datasets. Models offering good performance for Polish text retrieval include:
@@ 6.1.   Methodology
    Sparse-dense hybrids - The final step in our procedure was to build a hybrid system that combines results from dense and sparse models. Since these representations can capture different characteristics of the input data, merging results from two indexes can improve the overall performance of the text retrieval system (Chen et al., 2022; Luan et al., 2021). Naive aggregation methods involve computing new weights for query-document pairs as a weighted average of the scores returned by each index. Unfortunately, due to the different scales of the scores, it’s difficult to set such weights that would provide good zero-shot results for different datasets. More complex systems use two-stage retrieval, in which the first stage retrieves a number of relevant documents from each index, and then the documents are sorted by the reranker model. However, such systems are considerably slower than purely retrieval-based solutions.
@@ 9.   Bibliographical References
Appendix A Appendix: Detailed results

### 10 Mar 2024 v2

@@ 4.2.   Web datasets
The PIRB benchmark includes nine datasets crawled from Polish websites. Our goal was to enrich the benchmark with real-world data written originally in Polish. To accomplish this, we identified websites with separate Q&A sections, some containing questions and answers written directly by the site’s editors, and others allowing content creation by registered users. We selected a diverse set of data sources covering various domains such as medicine, law, business, physics, and linguistics. The resulting datasets consist of natural questions, formulated mostly as complete, grammatically correct sentences, as opposed to search engine-oriented datasets such as MS MARCO, in which queries are short and often consist of concatenated keywords. For each data source, we implemented a separate parser to enable precise extraction of question and answer content. We also conducted data cleaning and anonymization, removing personal and contact information. A summary of statistics regarding the collected datasets is presented in Table 1.
@@ 4.3.   GPT-exams
    In the last step, we manually removed from the dataset the cases in which the model refused to answer the question. We searched for phrases such as "model języka" (language model), "nie jestem" (I’m not), or "nie mogę" (I can’t). However, such cases were rare, we found less than 10 refusals for the entire dataset.
@@ 5.1.   Baseline methods
  • BM25 (Robertson et al., 2009) is a popular and effective term-based ranking function used in information retrieval since the 1980s. It is an extension of TF-IDF weighting scheme, providing a balanced approach to text ranking, considering both term frequency and document specificity. In our experiments, we used the implementation of BM25 available in anserini (Yang et al., 2018). We applied the default Polish analyzer, which performs the lemmatization of words with the Morfologik [^5] https://github.com/morfologik/morfologik-stemming library.
  • Sparse Lexical and Expansion (SPLADE) model (Formal et al., 2021b, a) is a family of modern term-based methods employing Transformer language models. In this approach, the masked language modeling (MLM) head is optimized to generate a vocabulary-sized weight vector adapted for text retrieval. SPLADE is a highly effective sparse retrieval ranking algorithm, achieving better performance than classic methods such as BM25. Unlike those methods, SPLADE directly uses subwords generated by the tokenizer as terms. We trained the Polish version following the same procedure as the SPLADE++ variant (Formal et al., 2022), utilizing hard negatives mined with an ensemble of cross-encoders (EnsembleDistil). We used Polish DistilRoBERTa (Dadas et al., 2020) as our base language model.
@@ 5.2.   Other methods
As part of our experiments, we evaluated over 20 publicly available Polish and multilingual dense text encoders, but only a few of them proved to be effective enough for practical use in retrieval tasks. In this publication, we present results only for the models that achieved an NDCG@10 score higher than the BM25 baseline. In the full ranking, which we have made available online666huggingface.co/spaces/sdadas/pirb, we include all of the evaluated retrievers, along with more detailed results covering individual datasets. Models offering good performance for Polish text retrieval include:
@@ 6.1.   Methodology
    Sparse-dense hybrids - The final step in our procedure was to build a hybrid system that combines results from dense and sparse models. Since these representations can capture different characteristics of the input data, merging results from two indexes can improve the overall performance of the text retrieval system (Chen et al., 2022; Luan et al., 2021). Naive aggregation methods involve computing new weights for query-document pairs as a weighted average of the scores returned by each index. Unfortunately, due to the different scales of the scores, it’s difficult to set such weights that would provide good zero-shot results for different datasets. More complex systems use two-stage retrieval, in which the first stage retrieves a number of relevant documents from each index, and then the documents are sorted by the reranker model. However, such systems are considerably slower than purely retrieval-based solutions since the most common reranker architectures require calculating the score for each query-document pair separately.
@@ 9.   Bibliographical References
Appendix A Appendix: Data pre-processing
This section describes our data pre-processing pipeline for Web Datasets. Since they contain questions and answers collected from the internet, additional cleaning and filtering steps are required to ensure the appropriate quality of evaluation data. For the GPT-exams, there is no need to perform the described steps because the model’s responses do not exhibit the same issues. The cleaning process primarily relies on heuristics and dictionary-based methods. More specifically, we employed the following steps:
  • Text normalization - We removed e-mails and www addresses using regular expressions. Additionally, special characters were removed, and sequences of whitespace characters were replaced with a single space.
  • Removing personal information - In order to anonymize the documents, we split the text into sentences and lemmatized each sentence. Then, we use a set of dictionaries of Polish first names, surnames, and words indicating contact information (such as "mail," "www," "tel", "address") to identify sentences containing such words and remove them from the texts.
  • Removing non-informative phrases - Using a dictionary-based method, we also removed phrases from the beginning and end of the text, both from questions and answers. We removed texts such as "good morning", "thank you in advance", "best regards", which do not contribute substantively to the content.
  • Numbering removal - In some sources, questions are numbered. For these datasets, we remove the number if it appears at the beginning of the question.
  • Removing questions with images - In two datasets (techpedia and onet), there may be questions that contain additional context in the form of an image. In the case of these datasets, we removed all questions containing phrases such as "picture", "photo", "image".
  • Removing short texts - Most erroneous and noisy samples come from short questions or answers. Moreover, short answers are often too general to be matched to a specific question without providing additional context. Therefore, we removed questions shorter than 10 characters and answers shorter than 50 characters (200 for abczdrowie and specprawnik, since those two datasets were of lower quality than the rest).
The source code of our preprocessing script, along with the dictionaries and execution arguments for each dataset, is available here: https://share.opi.org.pl/s/oMFJqJ5sSWjzFSX.
Appendix B Appendix: Error analysis
In this section, we present conclusions from error analysis based on the results obtained by our standalone retrieval models and hybrid methods. In order to further investigate the results of our methods, we conducted two comparisons. Below is the summary from the analysis.
Best of our dense retrievers (distilled and fine-tuned polish-roberta-large-v2) vs. previous state-of-the-art (multilingual-e5-large):
  • We divided all datasets into subgroups based on average query length: short (less than 10 words), medium (10-20 words), and long (more than 20 words). Our model outperforms E5 on 53% of short, 50% of medium, and 83% of long datasets.
  • PolEval-2022 is the only taks group, in which E5 performs significantly better. The group consists of seven datasets, all with a similar structure, comprising short natural queries and relatively short passages (below 50 words).
  • On the other hand, our model achieves better results on datasets from the BEIR-PL and Web Datasets groups, which consist of a more diverse sets of tasks. BEIR-PL also includes datasets with short queries from search engines, such as MS MARCO, NFCorpus, or DBPedia. On these datasets, our model has a few points of NDCG@10 advantage over E5.
  • E5 performs better or comparably to our model on trivia-like datasets, while our model exhibits better generalization ability to specialized domains.
Our standalone retrievers vs. hybrid retrieval:
  • The use of a hybrid retrieval improves results primarily for queries containing named entities or specialized terminology. Such cases can benefit from term-based matching of sparse indexes.
  • In addition, the use of the hybrid approach improves results for task groups, in which the standalone model performed worse or comparably to the original E5 model, particularly for PolEval-2022 and MAUPQA tasks. In the other groups, the differences are small.
Appendix C Appendix: Detailed results